OpenStudy (unklerhaukus):

\[\neg[\forall x[x < 0 \Rightarrow\exists y (y^2 = x)]]\iff\exists x[x\geq0\Rightarrow\forall y(y^2\neq x)]\]

OpenStudy (unklerhaukus):

am i right/

OpenStudy (unklerhaukus):

no wait , i see error

OpenStudy (unklerhaukus):

the negation of an implication

OpenStudy (helder_edwin):

the negation of \(\to\) is \(\wedge\)

OpenStudy (unklerhaukus):

yeah

OpenStudy (helder_edwin):

\[ \large \neg(p\to q)\equiv p\wedge\neg q \]

OpenStudy (unklerhaukus):

\[\exists x[x<0\wedge\forall y(y^2\neq x)]\]

OpenStudy (swissgirl):

Ya that makes more sense

OpenStudy (swissgirl):

hmmm you edited your question but its still incorrect

OpenStudy (unklerhaukus):

\[\neg\left[\forall x[x < 0 \Rightarrow\exists y (y^2 = x)]\right]\iff\exists x\left[x<0\land\forall y(y^2\neq x)\right]\]

OpenStudy (swissgirl):

Yes thats the answer

OpenStudy (unklerhaukus):

thanks to you both

OpenStudy (swissgirl):

not that I helped :P

OpenStudy (helder_edwin):

u r welcome